OpenStudy (anonymous):

evaluate 2_3^1*3_2^1 ,giving your answer as a mixed number..

OpenStudy (anonymous):

\(2_3^1\times3_2^1\)... I don't think that's what you mean.

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

We don't know what you're asking. Look at your post again.

OpenStudy (anonymous):

@badreferences yes! thats what my question mean.

OpenStudy (anonymous):

\(2\frac{1}{3}\cdot3\frac{1}{2}\neq2_3^1\cdot3_2^1\), be careful of what you ask.

OpenStudy (anonymous):

Also, the answer is \(\frac{7}{3}\cdot\frac{7}{2}=\frac{49}{6}\).

OpenStudy (anonymous):

Woops, yeah, I shouldn'tve given the answer immediately.

OpenStudy (anonymous):

no, just kidding but I wasn't thinking!

OpenStudy (anonymous):

Here's that youtube vid anyway. If she wants to know HOW to do this, then she'll really get a lot out of this. http://www.youtube.com/watch?v=aL3znpsf8X0